# Non-relativistic transport from frame-indifferent kinetic theory

### v1

@@ Acknowledgements
R.B. and P.S. have been supported, in part, by the Polish National Science Centre (NCN) Sonata Bis Grant 2019/34/E/ST3/00405. F.P.-B. has received funding from the Norwegian Financial Mechanism 2014-2021 via the NCN, POLS Grant 2020/37/K/ST3/03390. P.M. has been supported through the International Max Planck Research School (IMPRS) for Quantum Dynamics and Control hosted at the Max Planck Institute for the Physics of Complex Systems.
@@ Appendix A Gradient expansion
We start by noticing that the following relation holds between the $k$-th and the $(k-1)$-th moment of the distribution function as defined in Eq. (71):
@@
In other words, the temporal component of the $k$-th moment is equal to the $(k-1)$-th moment. Therefore, we can concentrate on finding the spatial components of the moments. We denote the purely spatial parts of $\mathcal{I}^{\alpha_{1}\alpha_{2}\cdots\alpha_{k}}$, $\mathcal{K}^{\alpha_{1}\alpha_{2}\cdots\alpha_{k}}$ and $\mathcal{S}^{\alpha_{1}\alpha_{2}\cdots\alpha_{k}}$ using a plain font:
@@
Note that the spatial projection used here is the one defined in Eq. (41) using $u^{\mu}$. Applying the projection operators $\bar{P}^{\alpha_{1}}_{\alpha_{1}^{\prime}}\bar{P}^{\alpha_{2}}_{\alpha_{2}^{\prime}}\cdots\bar{P}^{\alpha_{k}}_{\alpha_{k}^{\prime}}$ to both sides of Eq. (73) in the torsionless case $H_{\mu\nu}=0$ gives
@@
The left-hand side of Eq. (128) can be expressed in terms of the spatial tensors $I^{\alpha_{1}\alpha_{2}\cdots\alpha_{k}}$. After some further manipulations described in detail in Appendix B.6, Eq. (128) takes the form
(£$\left(\pounds_{u}+\nabla_{\mu}u^{\mu}\right)I^{\alpha_{1}\cdots\alpha_{k}}$ $+\nabla_{\mu}I^{\mu\alpha_{1}\cdots\alpha_{k}}+\big{[}h^{\alpha_{k}\sigma}\left(u^{\mu}I^{\alpha_{1}\cdots\alpha_{k-1}}+2I^{\mu\alpha_{1}\cdots\alpha_{k-1}}\right)+\cdots$
@@
£$\pounds_{u}I^{\alpha_{1}\cdots\alpha_{k}}=u^{\mu}\partial_{\mu}I^{\alpha_{1}\cdots\alpha_{k}}-I^{\mu\alpha_{2}\cdots\alpha_{k}}\partial_{\mu}u^{\alpha_{1}}-\cdots-I^{\alpha_{1}\cdots\alpha_{k-1}\mu}\partial_{\mu}u^{\alpha_{k}}\,.$ (130)
@@
$\left[T_{(k)}I^{(k)}\right]^{\alpha_{1}\alpha_{2}\cdots\alpha_{k}}$ (£$\equiv\left(\pounds_{u}+\theta\right)I^{\alpha_{1}\alpha_{2}\cdots\alpha_{k}}+kI^{\mu(\alpha_{2}\cdots\alpha_{k}}\bar{f}^{\,\nu|\alpha_{1})}g_{\mu\nu}\,,$ (132)
@@
Within the relaxation time approximation described in Sec. VI, the general equation (131) can be further rewritten. Firstly, the equations for the moments $k=0$, $k=1$, and the trace of the equation for $k=2$ can be equivalently expressed as
@@
Equations (136)-(138) are in fact equal to equations (102). Secondly, the equations of motion for the out-of-equilibrium moments $\delta I^{\alpha\beta\cdots\gamma}$ take the form:
@@
where $S^{(k)}$ is a symmetric tensor defined using Eq. (127) and Eq. (106), with $k$ indices that we leave implicit, $\mathcal{T}_{(k)}$ is an operator that combines $T_{(k)}$ with the action of the collision operator
$\left[\mathcal{T}_{(k)}\delta I^{(k)}\right]^{\alpha_{1}\alpha_{2}\cdots\alpha_{k}}=\left(\pounds_{u}+\theta\right)\delta I^{\alpha_{1}\alpha_{2}\cdots\alpha_{k}}+k\delta I^{\mu(\alpha_{2}\cdots\alpha_{k}}\bar{f}^{\nu|\alpha_{1})}g_{\mu\nu}+\frac{1}{\tau_{\mathrm{coll}}}\delta I^{\alpha_{1}\alpha_{2}\cdots\alpha_{k}}\,,$ (141)
and we denote the traceless symmetric part of a tensor by the angle brackets:
We will now examine requirements for the hydrodynamic approach to be valid. As discussed in the main text, the condition is that the frequencies of perturbations around equilibrium are much smaller than the relaxation rate $1/\tau_{\mathrm{coll}}$. Formally, we can define the characteristic frequency of the perturbation $\omega$ as the maximal magnitude of the eigenvalue of the operator £$\pounds_{u}$ acting on $\delta\mathcal{I}^{(k)}$, and demand $\omega\ll 1/\tau_{\mathrm{coll}}$. Furthermore, as seen in Eq. (63), in equilibrium $\theta$ and $\sigma^{\mu\nu}$ are both zero. We will demand that the non-equilibrium contributions to $\mathcal{T}_{(k)}\delta I^{(k)}$, i.e., the terms containing $\theta$ and $\sigma^{\mu\nu}$, are subleading with respect to $\frac{1}{\tau_{\mathrm{coll}}}\delta I^{(k)}$. If we denote by $\sigma$ the maximal magnitude of the eigenvalue of the operator $\sigma^{\mu\alpha}g_{\mu\nu}$ acting on $\delta I^{(k)}$, the assumptions we make can be written as $\omega,\sigma,\theta\ll 1/\tau_{\mathrm{coll}}$. We can also estimate that $\delta I^{(k+1)}\sim u_{\mathrm{char}}\delta I^{(k)}$, where $u_{\mathrm{char}}$ is the magnitude of perturbation to velocity $u^{\mu}$. Then, from the definition of $\sigma$ and $\theta$ we can estimate $D_{(k)}\delta I^{(k+1)}\sim(\sigma\delta I^{(k)},\theta\delta I^{(k)})\ll 1/\tau_{\mathrm{coll}}\delta I^{(k)}$, so that the gradient term $D_{(k)}\delta I^{(k+1)}$ is also subleading. This forms the basis of the gradient expansion: the terms in Eq. (140) containing gradients are considered subleading and an expansion of the out-of-equilibrium moments $\delta I^{(k)}$ in the powers of $\tau_{\mathrm{coll}}\omega$, $\tau_{\mathrm{coll}}\sigma$ and $\tau_{\mathrm{coll}}\theta$ can be carried out.
@@
Analogically, from Eq. (140) we can deduce that up to the first order in gradients
@@
Following Eq. (144), the stress tensor $\tau^{\mu\nu}$ can be obtained by inverting the operators $\mathcal{T}_{(2)}$. In the hydrodynamic regime $\mathcal{T}_{(2)}$ can be treated as a four-rank tensor such that
@@
We will also assume that since vorticity $2\nabla^{[\mu}u^{\nu]}=f^{\mu\nu}$ can be non-zero in equilibrium, the vorticity scalar defined in Eq. (107) is a zeroth-order quantity that has to be taken account of. In two dimensions we will choose to do calculations in polar coordinates $(\rho,\phi)$. In the basis $\left\{\tau^{\rho\rho},\tau^{\rho\phi},\tau^{\phi\phi}\right\}$
@@
Inverting this matrix we obtain
@@
In three dimensions, we work in the cylindrical coordinate system $(z,\rho,\phi)$ where the $z$ direction is aligned with the vorticity vector defined in Eq. (115), i.e., $B_{\mu}dx^{\mu}=Bdz$. In the basis $\left\{\mathrm{tr}[\tau]/2,(\tau^{\rho\rho}-\rho^{2}\tau^{\phi\phi})/2,\tau^{\rho\phi},\tau^{\rho z},\tau^{\phi z},\tau^{\langle zz\rangle}\right\}$,
@@
Analogically, to calculate $q^{\mu}$, one needs to invert the operator $\mathcal{T}_{(3)}$, as seen from the definition $q^{\mu}=\frac{m}{2}g_{\alpha\beta}\delta I^{\alpha\beta\mu}$ and Eq. (146). In two dimensions, in the basis $\left\{\delta I^{\rho\rho\rho},\delta I^{\rho\rho\phi},\delta I^{\rho\phi\phi},\delta I^{\phi\phi\phi}\right\}$ it reads
@@
Inverting this matrix and contracting with $g_{\beta\gamma}$ one recovers Eq. (113). Similarly, in three dimensions, in the basis $\left\{\delta I^{\rho\rho\rho},\delta I^{\rho\rho\phi},\delta I^{\rho\phi\phi},\delta I^{\phi\phi\phi},\delta I^{\rho\rho z},\delta I^{\rho\phi z},\delta I^{\phi\phi z},\delta I^{\rho zz},\delta I^{\phi zz},\delta I^{zzz}\right\}$
@@
Inverting this matrix and contracting with $g_{\beta\gamma}$ one recovers Eq. (121).
@@ B.1 Noether charge
Under an arbitrary infinitesimal coordinate reparametrization $x^{\mu}\rightarrow x^{\mu}+\xi^{\mu}$, the on-shell variation of the Lagrangian is
$\delta_{\mathrm{on}}\mathcal{L}=\xi^{\mu}\frac{\partial\mathcal{L}}{\partial x^{\mu}}+\dot{\xi}^{\mu}\frac{\partial\mathcal{L}}{\partial\dot{x}^{\mu}}=\frac{d}{d\lambda}\left(\xi^{\mu}\frac{\partial\mathcal{L}}{\partial\dot{x}^{\mu}}\right)\,,$ (156)
where we have used the Euler-Lagrange equations to get the last equality. On the other hand, the Lagrangian for a point particle in NC geometry is given by Eq. (21) as
@@
where $N=\tau_{\rho}\dot{x}^{\rho}$. The variation of $\mathcal{L}$ under the reparametrization $x^{\mu}\rightarrow x^{\mu}+\xi^{\mu}$ can be calculated explicitly as
$\delta\mathcal{L}=$ $\frac{m}{N}h_{\alpha\beta}\dot{x}^{\alpha}\dot{x}^{\rho}\partial_{\rho}\xi^{\beta}+\frac{m}{2N}\dot{x}^{\alpha}\dot{x}^{\beta}\xi^{\mu}\partial_{\mu}h_{\alpha\beta}-\frac{m}{2N^{2}}h_{\alpha\beta}\dot{x}^{\alpha}\dot{x}^{\beta}\left(\tau_{\mu}\dot{x}^{\rho}\partial_{\rho}\xi^{\mu}+\dot{x}^{\mu}\xi^{\rho}\partial_{\rho}\tau_{\mu}\right)$
$+mm_{\alpha}\dot{x}^{\rho}\partial_{\rho}\xi^{\alpha}+m\dot{x}^{\alpha}\xi^{\mu}\partial_{\mu}m_{\alpha}$
$=$ $\frac{m}{2N}\dot{x}^{\alpha}\dot{x}^{\beta}\left(\xi^{\mu}\partial_{\mu}h_{\alpha\beta}+2h_{\alpha\mu}\partial_{\beta}\xi^{\mu}\right)-\frac{m}{2N^{2}}h_{\alpha\beta}\dot{x}^{\alpha}\dot{x}^{\beta}\dot{x}^{\mu}\left(\xi^{\rho}\partial_{\rho}\tau_{\mu}+\tau_{\rho}\partial_{\mu}\xi^{\rho}\right)$
$+m\dot{x}^{\alpha}\left(\xi^{\mu}\partial_{\mu}m_{\alpha}+m_{\mu}\partial_{\alpha}\xi^{\mu}\right)$
$=$ $\frac{m}{2N}\dot{x}^{\alpha}\dot{x}^{\beta}\pounds_{\xi}h_{\alpha\beta}-\frac{m}{2N^{2}}h_{\alpha\beta}\dot{x}^{\alpha}\dot{x}^{\beta}\dot{x}^{\mu}\pounds_{\xi}\tau_{\mu}+m\dot{x}^{\alpha}\pounds_{\xi}m_{\alpha}\,.$ (158)
Let us denote all the transformations as $\chi=\left(\xi^{\mu},\psi_{\mu},\Lambda\right)$, where $\xi^{\mu}$ is an infinitesimal coordinate reparameterization, $\psi_{\mu}$ is a Milne boost, and $\Lambda$ is a $U(1)$ gauge transformation. If the parameters satisfy the Killing conditions in Eq. (30), the transformation $\chi$ leaves the action invariant up to a boundary term, and the variation of the Lagrangian under $\chi$ is called the symmetry variation [63, 64, 65]. If we denote by $\chi_{K}$ the set of Killing parameters, then we have
@@
£$\Rightarrow\pounds_{\xi_{K}}h_{\alpha\beta}=\psi^{K}_{\alpha}\tau_{\beta}+\psi^{K}_{\beta}\tau_{\alpha}\,,$ (160)
@@
£$\Rightarrow\pounds_{\xi_{K}}m_{\alpha}=-\psi^{K}_{\alpha}-\partial_{\alpha}\Lambda_{K}\,.$ (161)
For a symmetry variation we can use Eqs. (159), (160) and (161) to simplify Eq. (158). Thus, the symmetry variation of $\mathcal{L}$ reads
$\delta_{\mathrm{sy}}\mathcal{L}=-m\frac{d}{d\lambda}\Lambda_{K}\,.$ (162)
If the arbitrary coordinate reparametrization $\xi^{\mu}$ in the on-shell variation in Eq. (156) is replaced by the Killing transformation $\xi^{\mu}=\xi^{\mu}_{K}$, then the left-hand sides of both the transformations are equal [62]. Equating the on-shell and the symmetry variation in Eqs. (156) and (162) we get
@@
where $\xi^{\mu}$ is a vector, $\pi_{\mu}$ is the momentum and $\Lambda$ is a gauge parameter. We define
@@
$\begin{split}\Gamma^{\mu}_{\alpha\beta}=&\,\hat{v}^{\mu}\partial_{\beta}\tau_{\alpha}+\frac{1}{2}h^{\mu\sigma}\left(\partial_{\alpha}\bar{h}_{\beta\sigma}+\partial_{\beta}\bar{h}_{\alpha\sigma}-\partial_{\sigma}\bar{h}_{\alpha\beta}\right)\\ &+\frac{1}{2}h^{\mu\sigma}\left(m_{\sigma}H_{\beta\alpha}+m_{\alpha}H_{\sigma\beta}+m_{\beta}H_{\sigma\alpha}\right)\,,\\ \hat{v}^{\mu}\bar{h}_{\mu\nu}=&\left(2v^{\sigma}m_{\sigma}-h^{\gamma\sigma}m_{\gamma}m_{\sigma}\right)\tau_{\nu}\,,\\ h^{\mu\sigma}\bar{h}_{\sigma\nu}=&\,\delta^{\mu}_{\nu}+\left(h^{\mu\sigma}m_{\sigma}-v^{\mu}\right)\tau_{\nu}\,.\end{split}$ (167)
@@
where $p^{\alpha}$ is the kinematic momentum satisfing the constraint $\tau_{\mu}p^{\mu}=m$. Acting with the Liouville’s operator on the Noether charge gives
@@
Using the identities in Eq. (167) as well as $\tau_{\rho}\left(\frac{p^{\rho}\tau_{\mu}}{m}-\delta^{\rho}_{\mu}\right)=0$ we have explicitly
$\begin{split}&\Gamma^{\gamma}_{\alpha\beta}\bar{h}_{\gamma\rho}\frac{p^{\rho}\tau_{\mu}}{m}-\Gamma^{\gamma}_{\alpha\beta}\bar{h}_{\mu\gamma}=\Gamma^{\gamma}_{\alpha\beta}\bar{h}_{\gamma\rho}\left(\frac{p^{\rho}\tau_{\mu}}{m}-\delta^{\rho}_{\mu}\right)\\ &=\frac{1}{2}\left[\left(\partial_{\alpha}\bar{h}_{\beta\rho}+\partial_{\beta}\bar{h}_{\alpha\rho}-\partial_{\rho}\bar{h}_{\alpha\beta}\right)+\left(m_{\alpha}H_{\rho\beta}+m_{\beta}H_{\rho\alpha}+m_{\rho}H_{\beta\alpha}\right)\right]\left(\frac{p^{\rho}\tau_{\mu}}{m}-\delta^{\rho}_{\mu}\right)\,,\end{split}$ (171)
@@
Let us first focus on terms containing $H_{\mu\nu}$. After expanding $\left(\frac{p^{\rho}\tau_{\mu}}{m}-\delta^{\rho}_{\mu}\right)$ in Eq. (171) there are nine such terms. Four of them vanish by the virtue of the antisymmetricity of $H_{\mu\nu}$, meaning that $p^{\alpha}p^{\beta}H_{\beta\alpha}=p^{\alpha}p^{\rho}H_{\rho\alpha}=p^{\rho}p^{\beta}H_{\rho\beta}=0$, and the remaining five mutually cancel, so that in the end the tensor $H_{\mu\nu}$ disappears from Eq. (170). After expanding the terms containing derivatives of $\bar{h}_{\mu\nu}$ most of them cancel out as well. In the end, Eq. (170) simplifies to

### v2

@@ Acknowledgements
Special thanks to Piotr Witkowski for his contributions and involvement in the early stages of this work. R.B. and P.S. have been supported, in part, by the Polish National Science Centre (NCN) Sonata Bis Grant 2019/34/E/ST3/00405. F.P.-B. has received funding from the Norwegian Financial Mechanism 2014-2021 via the NCN, POLS Grant 2020/37/K/ST3/03390. P.M. has been supported through the International Max Planck Research School (IMPRS) for Quantum Dynamics and Control hosted at the Max Planck Institute for the Physics of Complex Systems.
@@ Appendix A Gradient expansion
By examining Eq. (63), the following relation between the $k$-th and the (k1)(k-1)-th moment of the distribution function can be derived:
@@
In other words, the temporal component of the $k$-th moment is equal to the (k1)(k-1)-th moment. Therefore, we can focus on finding only the spatial components of the moments. We distinguish the purely spatial parts of $\mathcal{I}^{\alpha_{1}\alpha_{2}\cdots\alpha_{k}}$, $\mathcal{K}^{\alpha_{1}\alpha_{2}\cdots\alpha_{k}}$ and $\mathcal{S}^{\alpha_{1}\alpha_{2}\cdots\alpha_{k}}$ by a plain font:
@@
Note that the spatial projection used here is the one defined in Eq. (25) using $u^{\mu}$. Applying the projection operators $\bar{P}^{\alpha_{1}}_{\alpha_{1}^{\prime}}\bar{P}^{\alpha_{2}}_{\alpha_{2}^{\prime}}\cdots\bar{P}^{\alpha_{k}}_{\alpha_{k}^{\prime}}$ to both sides of Eq. (65) gives, in the torsionless case $H_{\mu\nu}=0$,
@@
The projection operators on the left-hand side of Eq. (131) can be shifted inside the covariant derivative at the price of generating new terms in the equation, which is described in detail in Appendix B.6. In the end, Eq. (131) takes the form
$\left(\pounds_{u}+\nabla_{\mu}u^{\mu}\right)I^{\alpha_{1}\cdots\alpha_{k}}$ $+\nabla_{\mu}I^{\mu\alpha_{1}\cdots\alpha_{k}}+\big[h^{\alpha_{k}\sigma}\left(u^{\mu}I^{\alpha_{1}\cdots\alpha_{k-1}}+2I^{\mu\alpha_{1}\cdots\alpha_{k-1}}\right)+\cdots$
@@
$\pounds_{u}I^{\alpha_{1}\cdots\alpha_{k}}=u^{\mu}\partial_{\mu}I^{\alpha_{1}\cdots\alpha_{k}}-I^{\mu\alpha_{2}\cdots\alpha_{k}}\partial_{\mu}u^{\alpha_{1}}-\cdots-I^{\alpha_{1}\cdots\alpha_{k-1}\mu}\partial_{\mu}u^{\alpha_{k}}\,.$ (133)
@@
$\left[T_{(k)}I^{(k)}\right]^{\alpha_{1}\alpha_{2}\cdots\alpha_{k}}$ $\equiv\left(\pounds_{u}+\theta\right)I^{\alpha_{1}\alpha_{2}\cdots\alpha_{k}}+kI^{\mu(\alpha_{2}\cdots\alpha_{k}}\bar{f}^{\,\nu|\alpha_{1})}g_{\mu\nu}\,$, (135)
@@
Within the relaxation time approximation, see Eq. (99), the general equation (134) can be further rewritten. Firstly, the equations for $k=0$, $k=1$, and the trace of the equation for $k=2$, can be expressed as
@@
$\frac{d}{2}nu^{\mu}\partial_{\mu}T+nT\nabla_{\mu}u^{\mu}+\tau^{\mu\nu}g_{\rho\mu}\nabla_{\nu}u^{\rho}+\nabla_{\mu}q^{\mu}=0\,$, (141)
i.e., they take the form of equations (103)-(105). Secondly, the equations of motion for the out-of-equilibrium moments $\delta I^{\alpha\beta\cdots\gamma}$ take the form:
@@
where $S^{(k)}$ is the spatial part of the operator defined in Eq. (108) with $k$ indices that we leave implicit, and $\mathcal{T}_{(k)}$ is a combination of $T_{(k)}$ and the collision operator:
$\left[\mathcal{T}_{(k)}\delta I^{(k)}\right]^{\alpha_{1}\alpha_{2}\cdots\alpha_{k}}=\left(\pounds_{u}+\theta\right)\delta I^{\alpha_{1}\alpha_{2}\cdots\alpha_{k}}+k\delta I^{\mu(\alpha_{2}\cdots\alpha_{k}}\bar{f}^{\nu|\alpha_{1})}g_{\mu\nu}+\frac{1}{\tau_{\mathrm{coll}}}\delta I^{\alpha_{1}\alpha_{2}\cdots\alpha_{k}}\,.$ (144)
We will now examine the requirements for the hydrodynamic approach to be valid. As discussed in the main text, the condition is that the frequencies of out-of-equilibrium perturbations are much smaller than the relaxation rate $1/\tau_{\mathrm{coll}}$. What complicates the matter is the fact that there are multiple frequency scales that can be defined in this context. To find them, we first introduce a set of operators: $\{\pounds_{u},\sigma^{\mu\alpha}g_{\mu\nu},\theta,\frac{T}{m}h^{\mu\nu}\nabla_{\mu}\nabla_{\nu}\}$. Then, we can formally define the set of characteristic frequencies of the out-of-equilibrium perturbations $\{\omega,\sigma,\theta,\omega_{T}^{2}\}$ as the maximal magnitudes of the eigenvalues of the above-defined operators acting on $\delta\mathcal{I}^{(k)}$. In the hydrodynamic regime, we require $\mathrm{max}\{\omega,\sigma,\theta,\omega_{T}\}\ll 1/\tau_{\mathrm{coll}}$. This brings about two simplifications: one of them is that the terms containing $\pounds_{u}$, $\theta$ and $\sigma^{\mu\nu}$ in Eq. (144) are subleading with respect to $\frac{1}{\tau_{\mathrm{coll}}}\delta I^{(k)}$; the other one is that on dimensional grounds we can estimate $\delta I^{(k+1)}\sim\mathrm{max}\{u_{\mathrm{char}},\sqrt{T/m}\}\delta I^{(k)}$, where $u_{\mathrm{char}}$ stands for the magnitude of velocity perturbation out of equilibrium, so that $D_{(k)}\delta I^{(k+1)}\sim(\sigma\delta I^{(k)},\theta\delta I^{(k)},\omega_{T}\delta I^{(k)})\ll 1/\tau_{\mathrm{coll}}\delta I^{(k)}$, and consequently in Eq. (143) the gradient term $D_{(k)}\delta I^{(k+1)}$ is subleading. This is the essence of the gradient expansion: all the out-of-equilibrium terms are assumed to be subleading, such that an expansion in powers of $\tau_{\mathrm{coll}}\omega,\tau_{\mathrm{coll}}\sigma,\tau_{\mathrm{coll}}\theta,\tau_{\mathrm{coll}}\omega_{T}$ can be carried out.
@@
Analogously, from Eq. (143) we can deduce that up to the first order in gradients
@@
Following Eq. (146), the stress tensor $\tau^{\mu\nu}$ can be obtained by inverting the operators $\mathcal{T}_{(2)}$. In the hydrodynamic regime, $\mathcal{T}_{(2)}$ can be treated as a four-rank tensor such that
@@
Since vorticity $2\nabla^{[\mu}u^{\nu]}=f^{\mu\nu}$ can be non-zero in equilibrium, we take the vorticity scalar defined in Eq. (109) to be a zeroth-order quantity that has to be taken into account. In two dimensions, we will choose to do calculations in polar coordinates $(\rho,\phi)$. In the basis $\left\{\tau^{\rho\rho},\tau^{\rho\phi},\tau^{\phi\phi}\right\}$,
@@
Inverting this matrix, we obtain
@@
In three dimensions, we work in the cylindrical coordinate system $(z,\rho,\phi)$, where the $z$ direction is aligned with the vorticity vector defined in Eq. (118), i.e., $B_{\mu}dx^{\mu}=Bdz$. In the basis $\left\{\mathrm{tr}[\tau]/2,(\tau^{\rho\rho}-\rho^{2}\tau^{\phi\phi})/2,\tau^{\rho\phi},\tau^{\rho z},\tau^{\phi z},\tau^{\langle zz\rangle}\right\}$,
@@
Analogously, to calculate $q^{\mu}$, one needs to invert the operator $\mathcal{T}_{(3)}$, as seen from the definition $q^{\mu}=\frac{m}{2}g_{\alpha\beta}\delta I^{\alpha\beta\mu}$ and Eq. (148). In two dimensions, in the basis $\left\{\delta I^{\rho\rho\rho},\delta I^{\rho\rho\phi},\delta I^{\rho\phi\phi},\delta I^{\phi\phi\phi}\right\}$, $\mathcal{T}_{(3)}$ reads
@@
Inverting this matrix and contracting with $g_{\beta\gamma}$, one recovers Eq. (114). Similarly, in three dimensions, in the basis $\left\{\delta I^{\rho\rho\rho},\delta I^{\rho\rho\phi},\delta I^{\rho\phi\phi},\delta I^{\phi\phi\phi},\delta I^{\rho\rho z},\delta I^{\rho\phi z},\delta I^{\phi\phi z},\delta I^{\rho zz},\delta I^{\phi zz},\delta I^{zzz}\right\}$,
@@
Inverting this matrix and contracting with $g_{\beta\gamma}$, one recovers Eq. (124).
@@ B.1 Noether charge
Under an arbitrary infinitesimal path deformation $x^{\mu}\rightarrow x^{\mu}+\tilde{\xi}^{\mu}$, the on-shell variation of the Lagrangian is
$\delta_{\mathrm{on}}\mathcal{L}=\tilde{\xi}^{\mu}\frac{\partial\mathcal{L}}{\partial x^{\mu}}+\dot{\tilde{\xi}}^{\mu}\frac{\partial\mathcal{L}}{\partial\dot{x}^{\mu}}=\frac{d}{d\lambda}\left(\tilde{\xi}^{\mu}\frac{\partial\mathcal{L}}{\partial\dot{x}^{\mu}}\right)\,$, (158)
where we have used the Euler–Lagrange equations to get the last equality. On the other hand, the Lagrangian for a point particle in NC geometry is given by Eq. (36) as
@@
where $N=\tau_{\rho}\dot{x}^{\rho}$. The variation of $\mathcal{L}$ under the shift $x^{\mu}\rightarrow x^{\mu}+\tilde{\xi}^{\mu}$ can be calculated explicitly as
$\delta\mathcal{L}=$ $\frac{m}{N}h_{\alpha\beta}\dot{x}^{\alpha}\dot{x}^{\rho}\partial_{\rho}\tilde{\xi}^{\beta}+\frac{m}{2N}\dot{x}^{\alpha}\dot{x}^{\beta}\tilde{\xi}^{\mu}\partial_{\mu}h_{\alpha\beta}-\frac{m}{2N^{2}}h_{\alpha\beta}\dot{x}^{\alpha}\dot{x}^{\beta}\left(\tau_{\mu}\dot{x}^{\rho}\partial_{\rho}\tilde{\xi}^{\mu}+\dot{x}^{\mu}\tilde{\xi}^{\rho}\partial_{\rho}\tau_{\mu}\right)$
$+mm_{\alpha}\dot{x}^{\rho}\partial_{\rho}\tilde{\xi}^{\alpha}+m\dot{x}^{\alpha}\tilde{\xi}^{\mu}\partial_{\mu}m_{\alpha}$
$=$ $\frac{m}{2N}\dot{x}^{\alpha}\dot{x}^{\beta}\left(\tilde{\xi}^{\mu}\partial_{\mu}h_{\alpha\beta}+2h_{\alpha\mu}\partial_{\beta}\tilde{\xi}^{\mu}\right)-\frac{m}{2N^{2}}h_{\alpha\beta}\dot{x}^{\alpha}\dot{x}^{\beta}\dot{x}^{\mu}\left(\tilde{\xi}^{\rho}\partial_{\rho}\tau_{\mu}+\tau_{\rho}\partial_{\mu}\tilde{\xi}^{\rho}\right)$
$+m\dot{x}^{\alpha}\left(\tilde{\xi}^{\mu}\partial_{\mu}m_{\alpha}+m_{\mu}\partial_{\alpha}\tilde{\xi}^{\mu}\right)$
$=$ $\frac{m}{2N}\dot{x}^{\alpha}\dot{x}^{\beta}\pounds_{\tilde{\xi}}h_{\alpha\beta}-\frac{m}{2N^{2}}h_{\alpha\beta}\dot{x}^{\alpha}\dot{x}^{\beta}\dot{x}^{\mu}\pounds_{\tilde{\xi}}\tau_{\mu}+m\dot{x}^{\alpha}\pounds_{\tilde{\xi}}m_{\alpha}\,$, (160)
Let us denote all the transformations as $\chi=\left(\tilde{\xi}^{\mu},\psi_{\mu},\Lambda\right)$, where $\tilde{\xi}^{\mu}$ is an infinitesimal coordinate reparameterization, $\psi_{\mu}$ is a Milne boost, and $\Lambda$ is a $U(1)$ gauge transformation. If the parameters satisfy the Killing conditions in Eq. (45), the transformation $\chi$ leaves the action invariant up to a boundary term, and the variation of the Lagrangian under $\chi$ is called the symmetry variation [64, 65, 66, 67]. If we denote by $\chi_{K}$ the set of Killing parameters, then we have
@@
$\Rightarrow\pounds_{\xi_{K}}h_{\alpha\beta}=\psi^{K}_{\alpha}\tau_{\beta}+\psi^{K}_{\beta}\tau_{\alpha}\,$, (162)
@@
$\Rightarrow\pounds_{\xi_{K}}m_{\alpha}=-\psi^{K}_{\alpha}-\partial_{\alpha}\Lambda_{K}\,$, (163)
For a symmetry variation, we can use Eqs. (161), (162) and (163) to simplify Eq. (160). Thus, the symmetry variation of $\mathcal{L}$ reads
$\delta_{\mathrm{sy}}\mathcal{L}=-m\frac{d}{d\lambda}\Lambda_{K}\,$, (164)
If the arbitrary coordinate reparametrization $\tilde{\xi}^{\mu}$ in the on-shell variation in Eq. (158) is replaced by the Killing transformation $\xi^{\mu}=\xi^{\mu}_{K}$, then the left-hand sides of both the transformations are equal [63]. Equating the on-shell and the symmetry variation in Eqs. (158) and (164) we get
@@
where $\xi^{\mu}$ is a vector, $\pi_{\mu}$ is the canonical momentum, and $\Lambda$ is a gauge parameter. We define
@@
$\begin{split}\Gamma^{\mu}_{\alpha\beta}=&\,\hat{v}^{\mu}\partial_{\beta}\tau_{\alpha}+\frac{1}{2}h^{\mu\sigma}\left(\partial_{\alpha}\bar{h}_{\beta\sigma}+\partial_{\beta}\bar{h}_{\alpha\sigma}-\partial_{\sigma}\bar{h}_{\alpha\beta}\right)+\frac{1}{2}h^{\mu\sigma}\left(m_{\sigma}H_{\beta\alpha}+m_{\alpha}H_{\sigma\beta}+m_{\beta}H_{\sigma\alpha}\right)\,,\\ \hat{v}^{\mu}\bar{h}_{\mu\nu}=&\left(2v^{\sigma}m_{\sigma}-h^{\gamma\sigma}m_{\gamma}m_{\sigma}\right)\tau_{\nu}\,,\\ h^{\mu\sigma}\bar{h}_{\sigma\nu}=&\,\delta^{\mu}_{\nu}+\left(h^{\mu\sigma}m_{\sigma}-v^{\mu}\right)\tau_{\nu}\,.\end{split}$ (169)
@@
where $p^{\alpha}$ is the kinematic momentum satisfying the constraint $\tau_{\mu}p^{\mu}=m$. Acting with the Liouville operator on the collisional invariant gives
@@
Using the identities in Eq. (169), as well as $\tau_{\rho}\left(\frac{p^{\rho}\tau_{\mu}}{m}-\delta^{\rho}_{\mu}\right)=0$, we have explicitly
[…]
@@
Let us first focus on the terms containing $H_{\mu\nu}$. After expanding $\left(\frac{p^{\rho}\tau_{\mu}}{m}-\delta^{\rho}_{\mu}\right)$ in Eq. (173), there are nine such terms. Four of them vanish by the virtue of the antisymmetricity of $H_{\mu\nu}$, which implies $p^{\alpha}p^{\beta}H_{\beta\alpha}=p^{\alpha}p^{\rho}H_{\rho\alpha}=p^{\rho}p^{\beta}H_{\rho\beta}=0$, and the remaining five mutually cancel, so that in the end the tensor $H_{\mu\nu}$ disappears from Eq. (172). Furthermore, after expanding the terms containing derivatives of $\bar{h}_{\mu\nu}$, most of them cancel out as well. In the end, Eq. (172) simplifies to